OpenStudy (anonymous):

Which description best describes the solution to the following system of equations? y = −one halfx + 9 y = x + 7 Lines y = −one halfx + 9 and y = x + 7 intersect the x-axis. Lines y = −one halfx + 9 and y = x + 7 intersect the y-axis. Line y = −one halfx + 9 intersects the origin. Line y = −one halfx + 9 intersects line y = x + 7.

OpenStudy (anonymous):

@Thebadyboy

OpenStudy (zale101):

Do the two function have the same y-intercept? If so, then they intersect ion is at the y-axis, if not then they don't intersect at the y-axis at all. Do the two function have the same x-intercept? If so, then they intersection is at the x-axis, if not then they don't intersect at the x-axis at all.

OpenStudy (anonymous):

i don't know @Zale101

OpenStudy (zale101):

If you have no idea what an origin mean, then take a look at this graph. Do you think the two lines are located at the origin (0,0)? dw:1433462825582:dw|

OpenStudy (anonymous):

no

OpenStudy (zale101):



OpenStudy (anonymous):

so would it be C?

OpenStudy (anonymous):

yes or no @Zale101

OpenStudy (zale101):

\(\Large y = −\frac{1}{2}x + 9\) \(\Large y = x + 7\) Does the functions intersect at (0,0) meaning do both of the equations has the input x=0 and y=0 as the output? Plug x=0 and see what you get for y in both of the equations. Can you do that?

OpenStudy (anonymous):

ugh this is so confusing

OpenStudy (zale101):

\(\large y = −\frac{1}{2}x + 9 \) what is y when x=0? \(\large y = x + 7\) what is y when x=0?

OpenStudy (anonymous):

8.5?!

OpenStudy (zale101):

\(\Large y = −\frac{1}{2}(0) + 9=?\) \( \large y = (0) + 7=?\)

OpenStudy (zale101):

8.5 is incorrect

OpenStudy (anonymous):

8.5 and 7

OpenStudy (zale101):

7 is correct for the second equation, but how did you get 8.5 for the first equation?

OpenStudy (anonymous):

-1/2 +9

OpenStudy (zale101):

-1/2 times 0 is not -1/2 it is 0

OpenStudy (anonymous):

oh so its 9 and 7

OpenStudy (zale101):

Yes

OpenStudy (anonymous):

9 intercepts the x axis and 7 intercepts the y axis

OpenStudy (anonymous):

correct?

OpenStudy (zale101):

so both of the equations have different y-intercept and it doesn't lie on the origin. None of them had y=0 when x=0, and they each had a different y-intercept. The first equation had (0,9) and the second equation had (0,7).

OpenStudy (anonymous):

would it be D??

OpenStudy (anonymous):

i think so